# Rethinking the Embodied Gap in Vision-and-Language Navigation: A Holistic Study of Physical and Visual Disparities

### v1

@@ Abstract
Recent Vision-and-Language Navigation (VLN) advancements are promising, but their idealized assumptions about robot movement and control fail to reflect physically embodied deployment challenges. To bridge this gap, we introduce VLN-PE, a physically realistic VLN platform supporting humanoid, quadruped, and wheeled robots. For the first time, we systematically evaluate several ego-centric VLN methods in physical robotic settings across different technical pipelines, including classification models for single-step discrete action prediction, a diffusion model for dense waypoint prediction, and a train-free, map-based large language model (LLM) integrated with path planning. Our results reveal significant performance degradation due to limited robot observation space, environmental lighting variations, and physical challenges like collisions and falls. This also exposes locomotion constraints for legged robots in complex environments. VLN-PE is highly extensible, allowing seamless integration of new scenes beyond MP3D, thereby enabling more comprehensive VLN evaluation. Despite the weak generalization of current models in physical deployment, VLN-PE provides a new pathway for improving cross-embodiment’s overall adaptability. We hope our findings and tools inspire the community to rethink VLN limitations and advance robust, practical VLN models. The code is available at here.
@@ 4 Baselines
Improved VLMaps. Train-free, map-based models have demonstrated significant potential, driven by advancements in vision-and-language models with open-set vocabulary capabilities and the versatility of LLMs. A representative work, VLMaps [21] leverages LLMs to parse complex natural language commands into a sequence of code-like subgoals that include object and positional relations (e.g., “robot.move_in_between(‘sofa’, ‘chair’)"). These subgoals are then localized on a semantic map. VLMaps employs LSeg [27] to ground 3D point cloud voxels into semantic embeddings, aligning them with text embeddings of subgoals via contrastive learning. The LLM generates executable Python code, enabling the robot to define functions, logical structures, and parameterized API calls. In our experiments, we observed that some landmarks referenced in VLN instructions may not be visible initially. To address this, we incorporate an exploration policy VLFM [62] for frontier detection. Specifically, when the robot fails to detect the target object from the current pixel embeddings, it performs a turn-around maneuver, moves to each frontier, and evaluates whether the viewpoint is likely to contain the next landmark using the image and text encoders from LSeg. Please refer to our appendix for more details.
@@ 5.1 Evaluation of VLN-CE Solutions in VLN-PE
Table 2: Evaluation of end-to-end methods on the GRU-VLN10 dataset using the Humanoid robot to assess model generalization in an out-of-MP3D-domain setting. ZS means the zero-shot implementation. “w/o FT" means without fine-tuning in the specific scenes.
@@
Zero-shot Performance. As shown in Tab. 1 and Fig. 4, models transferred directly from VLN-CE to VLN-PE experience significant performance drops. Specifically, Seq2Seq-Full (#$\#2$), CMA-Full (#$\#3$), and NaVid (#$\#6$) see SR declines of 10%, 16%, and 18%, respectively. Generally, NaVid exhibits better generalization compared to smaller models. As its training code is unavailable, we assess its zero-shot performance as a 7B-parameter model. Despite some performance drop, NaVid achieves the highest zero-shot navigation results, highlighting the potential of MLLMs for VLN. Additionally, as a zero-shot, map-based LLM solution, VLMaps (#$\#12$) demonstrates reasonable results in the VLN-PE setting. This suggests that, beyond end-to-end action prediction, map-based intelligent navigation could also be a viable approach for VLN.
In-domain Training and Fine-tuning. Notably, Seq2Seq-Full and CMA-Full were trained on Habitat with extensive data augmentation, including dagger-based training and EnvDrop-augmented datasets [51], totaling 175K additional samples. Despite this, they underperform compared to models trained from scratch on in-domain data collected within VLN-PE—without any augmentation (#$\#7$ and #$\#8$). Interestingly, fine-tuning Seq2Seq and CMA (denoted as Seq2Seq+ and CMA+) with their SoTA weights using the training dataset from VLN-PE significantly improves performance. In particular, CMA+ (#$\#11$) surpasses NaVid’s zero shot performance, achieving SR 28.72 vs. 21.58 and SPL 24.24 vs. 17.45 on val-seen. These results indicate that existing VLN models tend to overfit to specific simulation platforms, resulting in poor generalization when directly transferred to new environments or settings. However, incorporating diverse domain data can further enhance overall navigation performance. Without additional augmentation, #$\#5$ explores the impact of observation height, aligning the agent’s camera height with H1, which improves transferability compared to #$\#4$ with the default height of 1.2m. Comparing #$\#9$ and #$\#7,\#8$, we find that RDP outperforms CMA and Seq2Seq when trained from scratch. This marks the first application of continuous dense low-level offset prediction in VLN, highlighting diffusion policy-based approaches as a promising research direction.
@@
Out-of-MP3D-domain Study. Since most VLN methods have been developed and evaluated within MP3D environments, we introduce 10 high-quality indoor scenes from GRUScenes [54] to expand the evaluation scope. To address the limited vocabulary size in the original CMA model, we replace its instruction encoder with Long-CLIP [63], denoted as CMA-CLIP. As shown in Tab. 2, fine-tuning small models (#$\#3$ and #$\#5$) using just 441 episodes from 3 scene leads to a significant performance boost compared to NaVid’s zero-shot results (e.g., RDP improves SR by 18.86% and SPL by 17.66% on average). Surprisingly, on the 3DGS-Lab-VLN dataset (Tab. 3), NaVid completely fails, continuously rotating and achieving only 5.81 SR. We suspect this is due to rendering noise introduced by 3DGS, which, while imperceptible to humans, may significantly disrupt the large model’s RGB perception. It is important to note that these results do not suggest that small models are sufficient, but rather highlight the substantial room for improvement in current MLLM-based navigation models.
@@ 5.2 Impact of Physical Controller
Collect w/ loco Eval w/ loco Val Seen Val Unseen
@@ 5.3 Impact of Cross-embodiment Data
Fig. 5 evaluates NaVid’s zero-shot performance on different robot types, revealing varying degrees of performance degradation. A key advantage of VLN-PE is its ability to seamlessly support diverse robot models for navigation and data collection. This raises our interest: Can cross-robot data improve model training, enabling a unified model that generalizes across different robot embodiments (“One-for-All")? To explore this, we collected R2R data from humanoid, quadrupedal, and wheeled robots within VLN-PE. Tab. 5 presents key findings from this study.
@@ Acknowledgments
We sincerely thank all our collaborators for their valuable support and contributions. This paper is supported by the National Natural Science Foundation of China under Grants (624B2105, 62473295, 62333017).
@@ A.1 Recurrent Diffusion Policy for VLN
We employ a transformer-based diffusion module [48] with one encoder layer and three decoder layers. During training, the ground-truth trajectory coordinates $T\times(\Delta x,\Delta y,\Delta yaw)$ are perturbed with random noise, and the network is trained to predict and remove this noise. The iterative denoising process follows DDPM [18]. Additionally, we introduce a self-attention-based stop prediction head to determine the current stop progress (from 00 to $1$). The stop signal is triggered if: All predicted actions from the diffusion head are below the threshold 0.1, or the stop progress output exceeds 0.8. The output of the RDP is shown in Fig. 7. During navigation, RDP predicts 8 future trajectory waypoints and executes 4 steps per iteration.
@@ A.2 Improved VLMaps
In the VLN-PE, we apply additional techniques to reduce the fall rate and stuck rate. For example, we implement an A* algorithm as the local planner, assigning higher costs to dilated and unexplored areas. When executing commands like self.move_forward(1), the robot may collide with obstacles if not properly oriented. To address, we define a cost function to identify the optimal node $n^{*}$ from the robot’s perspective: $n^{*}=\mathop{\arg\min}\limits_{n}(||dist(n,x_{0})-dist(x_{g},x_{0})||+\alpha\gamma)$, where $x_{g}$ is the goal position, $x_{0}$ is the current position, $\alpha=0.25$ is a weight parameter, and $\gamma$ is the angle required to face the target. This ensures the robot slightly reorients itself before moving forward, minimizing collision risks. An example of the improved VLMaps is shown in Fig. 9. Compared to end-to-end methods, map-based modular approaches offer more explainable and reliable results. However, their performance heavily depends on mapping and localization accuracy, which could limit the practical deployment.
@@ A.3 Experimental Details
All training experiments are conducted using NVIDIA RTX 4090 GPUs. The CMA and Seq2Seq models are trained on a single GPU with a batch size of 2, requiring approximately one day to converge. The RDP model is trained on 4 GPUs using PyTorch’s DataParallel module, with a total batch size of 4, and completes training in around two days. All models are optimized using the AdamW optimizer with a learning rate of $1\times 10^{-4}$. The maximum trajectory length is set to 200. For evaluation, the CMA model requires approximately 4 hours to complete a full evaluation on the R2R-CE benchmark when run in parallel on 8 GPUs.
@@ A.5 Metrics
Metrics. Following standard VLN evaluation protocols [2, 25], we use five primary metrics: Trajectory Length (TL), measured in meters; Navigation Error (NE), which quantifies the distance between the predicted and actual stop locations; Success Rate (SR), indicating how often the predicted stop location falls within a predefined distance of the true location; Oracle Success Rate (OS), which assesses the frequency with which any point along the predicted path is within a certain distance of the goal; and Success Rate weighted by Inverse Path Length (SPL), which balances success rate with path efficiency. As physical realism is a key focus of this work, we introduce two more metrics: Fall Rate (FR), which measures the frequency of unintended falls, and Stuck Rate (StR), which quantifies instances where the agent becomes immobilized. Specifically, “Fall" is the robot having a roll $>$ 15° or pitch $>$ 35°, or a center-of-mass–to-foot height below a robot-specific threshold. “Stuck" is defined as both position and heading change $<$ 0.2m and 15° for 50 steps.
@@ A.7 Fine-tune on the specific datasets
To better evaluate out-of-MP3D-style domain generalization, we collect additional VLN datasets using GRUScenes and 3DGS-rendered environments. Since these datasets are primarily used for evaluation, only a small portion of the data is allocated for training, while the majority is reserved for testing. For CMA and RDP, all training experiments use a learning rate of 1e-4 with a cosine learning schedule. In Tab. 2 and Tab. 3, “w/o FT" refers to direct zero-shot transfer using VLN-PE-R2R-trained weights for evaluation without further in-domain fine-tuning in these new scenes. On the GRU-VLN10 dataset, small models significantly improved after 10 epochs of fine-tuning, whereas the SoTA large model NaVid showed limited zero-shot performance. This highlights the limited diversity of existing VLN benchmarks, which could not fully assess model generalization.

### v2

@@ Abstract
Recent Vision-and-Language Navigation (VLN) advancements are promising, but their idealized assumptions about robot movement and control fail to reflect physically embodied deployment challenges. To bridge this gap, we introduce VLN-PE, a physically realistic VLN platform supporting humanoid, quadruped, and wheeled robots. For the first time, we systematically evaluate several ego-centric VLN methods in physical robotic settings across different technical pipelines, including classification models for single-step discrete action prediction, a diffusion model for dense waypoint prediction, and a train-free, map-based large language model (LLM) integrated with path planning. Our results reveal significant performance degradation due to limited robot observation space, environmental lighting variations, and physical challenges like collisions and falls. This also exposes locomotion constraints for legged robots in complex environments. VLN-PE is highly extensible, allowing seamless integration of new scenes beyond MP3D, thereby enabling more comprehensive VLN evaluation. Despite the weak generalization of current models in physical deployment, VLN-PE provides a new pathway for improving cross-embodiment’s overall adaptability. We hope our findings and tools inspire the community to rethink VLN limitations and advance robust, practical VLN models. The code is available at https://crystalsixone.github.io/vln_pe.github.io.
@@ 4 Baselines
Improved VLMaps. Train-free, map-based models have demonstrated significant potential, driven by advancements in vision-and-language models with open-set vocabulary capabilities and the versatility of LLMs. A representative work, VLMaps [21] leverages LLMs to parse complex natural language commands into a sequence of code-like subgoals that include object and positional relations (e.g., “robot.move_in_between(‘sofa’, ‘chair’)”). These subgoals are then localized on a semantic map. VLMaps employs LSeg [27] to ground 3D point cloud voxels into semantic embeddings, aligning them with text embeddings of subgoals via contrastive learning. The LLM generates executable Python code, enabling the robot to define functions, logical structures, and parameterized API calls. In our experiments, we observed that some landmarks referenced in VLN instructions may not be visible initially. To address this, we incorporate an exploration policy VLFM [62] for frontier detection. Specifically, when the robot fails to detect the target object from the current pixel embeddings, it performs a turn-around maneuver, moves to each frontier, and evaluates whether the viewpoint is likely to contain the next landmark using the image and text encoders from LSeg. Please refer to our appendix for more details.
@@ 5.1 Evaluation of VLN-CE Solutions in VLN-PE
Table 2: Evaluation of end-to-end methods on the GRU-VLN10 dataset using the Humanoid robot to assess model generalization in an out-of-MP3D-domain setting. ZS means the zero-shot implementation. “w/o FT” means without fine-tuning in the specific scenes.
@@
Zero-shot Performance. As shown in Tab. 1 and Fig. 4, models transferred directly from VLN-CE to VLN-PE experience significant performance drops. Specifically, Seq2Seq-Full (#2\#2), CMA-Full (#3\#3), and NaVid (#6\#6) see SR declines of 10%, 16%, and 18%, respectively. Generally, NaVid exhibits better generalization compared to smaller models. As its training code is unavailable, we assess its zero-shot performance as a 7B-parameter model. Despite some performance drop, NaVid achieves the highest zero-shot navigation results, highlighting the potential of MLLMs for VLN. Additionally, as a zero-shot, map-based LLM solution, VLMaps (#12\#12) demonstrates reasonable results in the VLN-PE setting. This suggests that, beyond end-to-end action prediction, map-based intelligent navigation could also be a viable approach for VLN.
In-domain Training and Fine-tuning. Notably, Seq2Seq-Full and CMA-Full were trained on Habitat with extensive data augmentation, including dagger-based training and EnvDrop-augmented datasets [51], totaling 175K additional samples. Despite this, they underperform compared to models trained from scratch on in-domain data collected within VLN-PE—without any augmentation (#7\#7 and #8\#8). Interestingly, fine-tuning Seq2Seq and CMA (denoted as Seq2Seq+ and CMA+) with their SoTA weights using the training dataset from VLN-PE significantly improves performance. In particular, CMA+ (#11\#11) surpasses NaVid’s zero shot performance, achieving SR 28.72 vs. 21.58 and SPL 24.24 vs. 17.45 on val-seen. These results indicate that existing VLN models tend to overfit to specific simulation platforms, resulting in poor generalization when directly transferred to new environments or settings. However, incorporating diverse domain data can further enhance overall navigation performance. Without additional augmentation, #5\#5 explores the impact of observation height, aligning the agent’s camera height with H1, which improves transferability compared to #4\#4 with the default height of 1.2m. Comparing #9\#9 and #7,#8\#7,\#8, we find that RDP outperforms CMA and Seq2Seq when trained from scratch. This marks the first application of continuous dense low-level offset prediction in VLN, highlighting diffusion policy-based approaches as a promising research direction.
@@
Out-of-MP3D-domain Study. Since most VLN methods have been developed and evaluated within MP3D environments, we introduce 10 high-quality indoor scenes from GRUScenes [54] to expand the evaluation scope. To address the limited vocabulary size in the original CMA model, we replace its instruction encoder with Long-CLIP [63], denoted as CMA-CLIP. As shown in Tab. 2, fine-tuning small models (#3\#3 and #5\#5) using just 441 episodes from 3 scene leads to a significant performance boost compared to NaVid’s zero-shot results (e.g., RDP improves SR by 18.86% and SPL by 17.66% on average). Surprisingly, on the 3DGS-Lab-VLN dataset (Tab. 3), NaVid completely fails, continuously rotating and achieving only 5.81 SR. We suspect this is due to rendering noise introduced by 3DGS, which, while imperceptible to humans, may significantly disrupt the large model’s RGB perception. It is important to note that these results do not suggest that small models are sufficient, but rather highlight the substantial room for improvement in current MLLM-based navigation models.
@@ 5.2 Impact of Physical Controller
Collect
w/ loco
Eval
w/ loco
Val Seen Val Unseen
@@ 5.3 Impact of Cross-embodiment Data
Fig. 5 evaluates NaVid’s zero-shot performance on different robot types, revealing varying degrees of performance degradation. A key advantage of VLN-PE is its ability to seamlessly support diverse robot models for navigation and data collection. This raises our interest: Can cross-robot data improve model training, enabling a unified model that generalizes across different robot embodiments (“One-for-All”)? To explore this, we collected R2R data from humanoid, quadrupedal, and wheeled robots within VLN-PE. Tab. 5 presents key findings from this study.
@@ Acknowledgments
We sincerely thank all our collaborators for their valuable support and contributions. This paper is supported by the National Natural Science Foundation of China under Grants (624B2105, 62473295, 62233013, 62333017). This work is funded in part by the National Key R&D Program of China (2022ZD0160201), and Shanghai Artificial Intelligence Laboratory.
@@ A.1 Recurrent Diffusion Policy for VLN
We employ a transformer-based diffusion module [48] with one encoder layer and three decoder layers. During training, the ground-truth trajectory coordinates $T\times(\Delta x,\Delta y,\Delta yaw)$ are perturbed with random noise, and the network is trained to predict and remove this noise. The iterative denoising process follows DDPM [18]. Additionally, we introduce a self-attention-based stop prediction head to determine the current stop progress (from 0 to 11). The stop signal is triggered if: All predicted actions from the diffusion head are below the threshold 0.1, or the stop progress output exceeds 0.8. The output of the RDP is shown in Fig. 7. During navigation, RDP predicts 8 future trajectory waypoints and executes 4 steps per iteration.
@@ A.2 Improved VLMaps
[…]
@@ A.3 Experimental Details
All training experiments are conducted using NVIDIA RTX 4090 GPUs. The CMA and Seq2Seq models are trained on a single GPU with a batch size of 2, requiring approximately one day to converge. The RDP model is trained on 4 GPUs using PyTorch’s DataParallel module, with a total batch size of 8, and completes training in around two days. All models are optimized using the AdamW optimizer with a learning rate of $1\times 10^{-4}$. The maximum trajectory length is set to 200. For evaluation, the CMA model requires approximately 4 hours to complete a full evaluation on the R2R-CE benchmark when run in parallel on 8 GPUs.
@@ A.5 Metrics
Metrics. Following standard VLN evaluation protocols [2, 25], we use five primary metrics: Trajectory Length (TL), measured in meters; Navigation Error (NE), which quantifies the distance between the predicted and actual stop locations; Success Rate (SR), indicating how often the predicted stop location falls within a predefined distance of the true location; Oracle Success Rate (OS), which assesses the frequency with which any point along the predicted path is within a certain distance of the goal; and Success Rate weighted by Inverse Path Length (SPL), which balances success rate with path efficiency. As physical realism is a key focus of this work, we introduce two more metrics: Fall Rate (FR), which measures the frequency of unintended falls, and Stuck Rate (StR), which quantifies instances where the agent becomes immobilized. Specifically, “Fall” is the robot having a roll >> 15° or pitch >> 35°, or a center-of-mass–to-foot height below a robot-specific threshold. “Stuck” is defined as both position and heading change << 0.2m and 15° for 50 steps.
@@ A.7 Fine-tune on the specific datasets
To better evaluate out-of-MP3D-style domain generalization, we collect additional VLN datasets using GRUScenes and 3DGS-rendered environments. Since these datasets are primarily used for evaluation, only a small portion of the data is allocated for training, while the majority is reserved for testing. For CMA and RDP, all training experiments use a learning rate of 1e-4 with a cosine learning schedule. In Tab. 2 and Tab. 3, “w/o FT” refers to direct zero-shot transfer using VLN-PE-R2R-trained weights for evaluation without further in-domain fine-tuning in these new scenes. On the GRU-VLN10 dataset, small models significantly improved after 10 epochs of fine-tuning, whereas the SoTA large model NaVid showed limited zero-shot performance. This highlights the limited diversity of existing VLN benchmarks, which could not fully assess model generalization.